OpenStudy (anonymous):

What is the slope of the line with an x-intercept of -2 and passing through (1,-4)?

OpenStudy (anonymous):

if you go to this site it will do it for u https://campus.ctuonline.edu/Controls/#

OpenStudy (anonymous):

x-intercept mean (-2,0) and the slope is \[\frac{-4-0}{1-(-2)}=-\frac{4}{3}\]\]